OpenStudy (anonymous):

What is the slope of a line parallel to the line whose equation is 2x – 3y = 6? A. 1/5 B. 3/3 C. 2/3 D. 4/3

TheSmartOne (thesmartone):

The slope of parallel lines are always the same. :)

OpenStudy (alexandervonhumboldt2):

parallel lines have same slope

OpenStudy (anonymous):

@mathstudent55

TheSmartOne (thesmartone):

So first let's make your equation into y=mx + b form

OpenStudy (anonymous):

@pooja

OpenStudy (anonymous):

Right

OpenStudy (anonymous):

yall need some help?

TheSmartOne (thesmartone):

So we have 2x - 3y = 6 add 3y on both sides subtract 6 on both sides what do you get? :P

OpenStudy (anonymous):

Oh snap, I have a hard time converting standrad form into slope intercept form

OpenStudy (anonymous):

@TheSmartOne HELP me with this one question after this pls

TheSmartOne (thesmartone):

It's not that hard. :) You just add some stuff here, subtract some stuff there. Divide by a number here. It's all pretty simple :)

TheSmartOne (thesmartone):

So we have 2x - 3y = 6 First let's add 3y on both sides. So that way you don't have too much of trouble dealing with negative numbers. :)

OpenStudy (anonymous):

2x= 6 + 3y right?

TheSmartOne (thesmartone):

correct! now subtract 6 on both sides :D

OpenStudy (anonymous):

4x = 3y?

TheSmartOne (thesmartone):

nope, we can't subtract the 2x with it. 2x has a variable 'x' with it. But 6 doesn't so we can't subtract it :P

OpenStudy (anonymous):

Oh so would it be 6 - 2x = 3y?

TheSmartOne (thesmartone):

Nope, \(\sf\Large 2x = 3y + 6\) \(\sf\Large 2x -6 = 3y + 6 - 6\) \(\sf\Large 2x -6 = 3y\) And just so that way y will b in front of the equation :P \(\sf\Large 3y = 2x - 6\)

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Then I divide 3 on both sides right? @TheSmartOne

TheSmartOne (thesmartone):

Correct! :D

OpenStudy (anonymous):

Yes! :D Okay then I got y = 2/3x - 2 right?

OpenStudy (anonymous):

So then the answer would be C! 2/3!

TheSmartOne (thesmartone):

correct! Good job! :D

OpenStudy (anonymous):

Awesome! Now for the next 7 xD